∠A and ∠B are vertical angles. If m∠A=(x+25)∘ and m∠B=(7x+1)∘, then find the value of x.

Answer by stanbon(75887)

∠A and ∠B are vertical angles. If m∠A=(x+25)∘ and m∠B=(7x+1)∘, then find the value of x.
-------
Since they are vertical angles, they are equal.
-------
Equation:
x+25 = 7x+1
6x = 24
x = 4
--------------------------
Cheers,
Stan H.
=================